A 5.00 Liter ballon of gas at 298K is cooled to 273K. What is the new volume in liters?
1
Expert's answer
2016-04-15T08:35:05-0400
According to the gas laws:
V1/T1 = V2/T2
5.00 L/298 K = V2/273 K
V2 = (273 K* 5.00 L)/298 K = 4.58 L

Answer:
The new volume is 4.58 L